# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v2

@@ VI Conclusion
This paper developed a differentially private DMPC strategy for linear discrete-time systems with coupled global constraints. We showed that the DMPC method relying on the conventional distributed dual-gradient algorithm is susceptible to eavesdropping attacks. To address this issue, we incorporated a DP noise injection mechanism into the distributed dual-gradient algorithm, enabling privacy preservation while maintaining accurate optimization convergence. Furthermore, a practical implementation approach for DMPC was proposed, which guarantees the feasibility and stability of the closed-loop system. Simulation results validated the effectiveness of the developed privacy-preserving DMPC strategy. Future work will extend the differentially private framework for systems with uncertainties (e.g., robust and stochastic DMPCs).

### v3

@@ VI Conclusion
This paper developed a differentially private DMPC strategy for linear discrete-time systems with coupled global constraints. We incorporated a DP noise injection mechanism into the distributed dual-gradient algorithm, enabling privacy preservation while maintaining accurate optimization convergence. Furthermore, a practical implementation approach for DMPC was proposed, which guarantees the feasibility and stability of the closed-loop system. Simulation results validated the effectiveness of the developed privacy-preserving DMPC strategy. Future work will extend the differentially private framework to directed communication networks and systems with uncertainties (e.g., robust DMPCs), and will evaluate the proposed framework on practical applications.